| Products > Test Equipment |
| RIGOL MSO1104Z MSO problems |
| (1/3) > >> |
| Tangent_Tracker:
Does anyone else use this scope with the 16 channel MSO? We got ours new last year and very quickly had problems with the logic analyzer. Signals became unpredictable and it would fail to calibrate. After talking to Rigol, and sending the analyzer back, then the oscilloscope, the problem seemed to be fixed. I will say at this time we were told not, under any circumstances to remove the module when under power, as this can cause damage to the board! We had not actually done this, but it struck me that a piece of equipment in a lab environment like this is very poorly designed if you have to power cycle the scope every time you connect or disconnect a bit of equipment that handles low voltage logic, and is designed specifically for that instrument! Fast forward 6 months and it seems two of the channels have now started playing up and do not respond to inputs. Neither can you see them responding during calibration, but ironically everything calibrates ok... We have been incredibly careful and ensured the module has only ever been removed or replaced without power! Maybe we haven't waited long enough after switching power off I don't know, that would be ridiculous! Has anyone else encountered any issues with this scope and the MSO module? Thanks. |
| hexreader:
Well, I am a few months late replying :) but... I have had no problems with calibration, but I do find some logic analyser inputs to be intermittent. Current suspicion is that the connectors at each end of the ribbon cable mate poorly due to plastic covers prevent full insertion. Loosening the covers while fitted and allowing a little more plug to protrude seems to have cured the problem for now, but I suspect that I will end up sanding away some of the cover plastic to allow better mating. Will let you know if I have further troubles. If anyone wants to open the Logic Head, be aware that there are three torx screws holding the two halves together. They are hidden under labels, so your device will not look pretty once you have cut holes in the labels. Inside are 8 dual comparators with differential outputs (LMH7322), resistors and caps for signal conditioning and terminators and little else. EDIT: Starting to think that covers are only a problem at the probe end of the ribbon cable, which is the less troublesome end. Seems to be intermittent connection at the scope end, even with no cover fitted. Maybe I just have tarnished pins on the plugs??? Still investigating.... Still employing guesswork as main diagnostic tool :( |
| hexreader:
Zip file with more pictures attached at PM request from modder_mike :) ... as for connection problems, I have found that removing plastic covers from the ribbon cables solves all but one connection problem. The connector at one end still has a single pin issue, either where pin does not mate, or IDC connection is bad. Yet to determine which, as most of my test gear is in storage. |
| hexreader:
.... and also attached is spreadsheet of the connector pin functions. Don't take it too seriously, as a lot of guesswork was involved. The signal pins are probably correct though. Most of the pins around signal pins are ground, but I didn't bother to check fully |
| Howardlong:
Apologies for the necroposting. I have some information that may well be of use regarding certain individual channels intermittently failing. I've had the same thing, and earlier on in its life, it seemed to be corrected by cleaning the connectors with IPA and blowing some air into them to dislodge fluff etc. Much more recently, I've found that this action didn't have much effect. The nature of the fault is that over time various channels would come and go, almost as if there were some dry joints or one of those BGA ball problems. Some channels were more sticky than others. Adjusting the threshold level on the LA has no effect, but if you supply an appropriate square wave into the dodgy channels, such as from the internal signal generator if it's a DS1000Z-S model, you usually see a glitch pulse on the LA at transitions, but no square wave. Today I took mine apaaaaart, it'd come to the point where almost half of the channels had become unreliable. There's a very short 68 pin IDC cable assembly with an IDC header soldered onto the main board. I discovered if I applied light pressure to the top of the header, it had an effect on the LA channels. So I gently eased off the IDC ribbon header clip. After liberally applying IPA over the ribbon/blades, I individually firmly pushed each of the 68 ribbon wires onto each of the respective header blades with a 1.5mm flat bladed screwdriver I had to hand in an attempt to remake the cold weld. I replaced the top clip of the header, and this seems to have resolved it, for now at least: it's been on for an hour with no ill effects, typically I'd have seen three or four channels randomly come and go in that time. It may be that I'll need to revisit if this problem recurs, perhaps to make the header clip maintain more pressure to the ribbon cable. Edit: FWIW, over the years, I've pulled out and reinserted the RPL1116 cable while the scope's been on many times with no ill effects. Edit2: The LA Probe Cal also passes after the above intervention. |
| Navigation |
| Message Index |
| Next page |